OpenStudy (anonymous):

PLEASE PLEASE PLEASE IN THE NAME OF THE GODS HELP ME !!!!!!!!! What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 A.) −2 B.) −3 C.) x can be any number as there are infinitely many solutions to the system. D.) There is no x value as there is no solution to the system

OpenStudy (anonymous):

Im going to try this, I think. ) D(412) 3*708-2*412=1300 708+412=1120 2) C x can be any number(infinte solutons) 3) B (x=−3)

OpenStudy (jamesj):

You're the one doing the exam, right?

OpenStudy (anonymous):

UHM NO . this is just algebra 1 .

OpenStudy (anonymous):

Well set one equation to have x on one side and y on the other. then plug that x into the other equation

OpenStudy (anonymous):

Wait I messed up, hold on.

OpenStudy (anonymous):

okay sure (:

OpenStudy (anonymous):

Step 1. 3x-6y=3 Step 2. 3x=3+6y Step 3. x=1+2y (Here you divide each side by 3) Step 4. You know that x = 1+2y, so you can plug that into the x of the other equation Step 5. 7*(1+2y) - 5y = -11 Step 6. 7 + 14y -5y= -11 Step 7. 7 + 9y = -11 Step 8. 9y = -11 - 7 Step 9. 9y = -18 Step 10. y = -2

OpenStudy (anonymous):

Tada :]

OpenStudy (lilg132):

i just done all that but got the answer wrong too long to to do it again lol i used the method of elimination instead of substitution

OpenStudy (anonymous):

Thanks, its much appreciated but it turns out I already had the answer ): sorry, I didn't notice till, like, now.

OpenStudy (anonymous):

It's okay. Just remember you can substitute